OpenStudy (anonymous):

csctheta= -5/12, tantheta>0, find cos 2theta a)-4sqrt21/25 b)4sqrt21/25 c)17/25 d)-17/25

OpenStudy (john_es):

Maybe you can use the relation, \[\cos(2\theta)=1-\frac{2}{\csc^2(\theta)}\]

OpenStudy (john_es):

I think not, you mean cos^2theta or cos(2*theta)?

OpenStudy (anonymous):

cos2*theta

OpenStudy (john_es):

Ok, then the formula should be correct.